Home‎ > ‎Reports‎ > ‎Crew100A‎ > ‎

day10


Crew 100A Log Book for February 21, 2011

Photos of the day





Commander's Check-In Report

Ivo Ferreira Reporting

Crew Physical Status: Good

Time Departed/Returned from EVA: EVA #8, #9

Brief Narrative of Field Mission Results:

The morning was spent on the habitability experiments and ended with an early lunch to strech the afternoon.

range extender support for our mobile unit (car) was built, the goal was to place the high gain antennas higher to reduce diffraction of the wifi signal due to the sand and therefore enhance the connection There were two more EVAs involving 4 of the crew members to test the effect of stress induced by communication, this time with one walking mission and another with the ATV. Simultaneously, at the Habitat, the between the Habitat computer and Max (Rover).

Later on, while some of us rechecked the site of the earlier Tiger mission, the others tested the range extender by using Max and the parrot together. Although the strength of the connection definitely increased we still run into some problems with loss of signal with Max when its camera was working (maybe due bandwidth requirements).

The Sun was already low, but we had just received the feedback of the SALM team (20th February) regarding the parrot flybys we did yesterday. They had checked the movies and pictures and wanted us to recheck more specific sites. The slow winds and our adrenaline were the trigger and there we went for another flybys whose result we will send today.

Engineering/Hab Maintenance: N/A

Plans for Tomorrow:  Finish the habitability experiments (still missing the debriefing sessions), continue with more EVAs and hopefully sort out the problem with Max.

Reports scheduled for today: Engineering, Commander, Journalist, Science, EVA #8 and EVA#9

Inventory: None

Support Requested: N/A

Miscellaneous: N/A


Commander's Report

Ivo Ferreira Reporting

Today I noticed that we are becoming like a machine. A well-oiled machine where each part already knows what will be its role even before it is asked to do it. Let see... we are only 5 but we managed to: follow the habitability experiments (3 of us), do two more EVAs to study the effect of stress due to communication (involving 4 of us), improve our range extender with a support on the our mobile unit (2 of us), finish the Tiger mission with the laboratory analysis (1 of us), test the enhanced mobile range extender (4 of us) and, finally, do some more flybys over the SALM sites (2 of us).

In fact I am sure that if for us the 24 hours of the day are not enough we will work nights as well!

We were very excited to receive the feedback of our late-night flights yesterday (with the parrot hovering two interesting sites near the habitat) from the research team in La Reunion.

They actually managed to sustain the motion sickness, checked the movies made by the parrot yesterday and decided on some sites that they wanted to see in more detail! So, we did not rest until we could comply with their requests... It is really motivating to be able to reiterate our work here on Mars with someone on Earth's La Reunion Island. The internet service here on Mars is amazing these days.

Tonight we are fancying meat so it's time for Romain to give up his "precious" last beef portion. From now on we will all become "forced" vegetarians.


Engineering Report

Kévin Theatre Reporting

Generator/Electricity:

· Kitty (Diesel Generator) (Used? Yes/No): Yes

· Honey (Gas Generator) (Used? Yes/No): No

· Battery Charge Level: Checked by DG

o 10.59 V everstart

o 12.37 V armor plate

o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is good


Fuel Status:

· Diesel (Full/0.75/0.5/0.25/Empty): 0.76

· Propane Tank (Full/0.75/0.5/0.25/Empty): 0.54

· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 11 Gallons

Notes/Comments:


Oil:

· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG

· Honey Oil Quantity (# of Quarts in storage): Leave it to DG

· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench

· Car Oil Quantity (# of Quarts in storage):


Water Status:

· Outside Potable Water Tank Level (inches from bottom): 14

· Trailer Potable Water Tank Level (inches from bottom): Full

· Hab Potable Water Tank Level (inches from bottom):

· Potable Water Meter Reading: 423451


Notes/Comments:


GreenHab:

· Condition of plants in Tank 1: green water

· Condition of plants in Tank 2: green things

· Condition of plants in Tank 3: green things

· Greenhab notes/ comments: The level of water of the 3 tanks has reduced a little. I attached some picture as you had asked yesterday.


Observatory:

· Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

· Used (yes/no): No

· Oil Checked (yes/no): No

· Tire Status: checked

· Odometer at end of day: 123973

· Notes/Comments on Hab Car: just to supply in electricity the relay for the Rover during our first attempt to do mission SALM.


ATV 1 (Opportunity):

· Used: yes

· Oil Checked: yes

· Fuel Consumed: Replaced

· Tire Status: filled

· Comments:


ATV 2 (Spirit):

· Used: No

· Oil Checked: No

· Fuel Consumed:

· Tire Status: filled

· Comments: Still not fixed


ATV 3 (Viking I):

· Used: No

· Oil Checked: No

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


ATV 4 (Viking II):

· Used: yes

· Oil Checked: yes

· Fuel Consumed: Replaced

· Tire Status: Filled

· Comments:


Heating and Ventilation:

· Thermostat setting upstairs (Farenheit): 70

· Actual temperature upstairs (Farenheit): 70

· Thermostat setting downstairs (Farenheit): 60

· Actual temperature downstairs (Farenheit): 60


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering

1. Notes/Comments:


Science Report

Ayako Ono Reporting

1. This morning, we started with habitability experiment, and Audrey, Kévin and Romain performed with colors, fragrances and natural sound. You have already received the explanation of this experiment.

2. Then, Kévin worked with the extender for the antenna.

3. Ivo, Romain, Ayako worked on the EVA #8 and Ivo, Romain, Audrey worked on the EVA #9 and these were successfully done.

4. Mission SALM was also well done. Ivo and Romain worked to take photos and videos from the samples which were asked.


Request:

I’d like to ask to borrow color bars for camera from someone who has it. If there is it, would be very helpful to adjust exact color for the habitability experiment. To take the picture of the lighting environment, I just need the color bar only once, before I leave from the MDRS.


EVA 8 Report

Romain Benchenafi Reporting

EVA #8, 14:42 - 15:07

Crew members : ONO Ayako, BENCHENAFI Romain

Site location : UTM 12S 0518331 4250490,  UTM 12S 0518331 4250395

Transit mode : Foot

Objective : Monitor factors of stress during an EVA

Result : Succeeded to comply with the two scientific objectives and with the emergency.

Lesson learned : The priority is to ensure the security of the crew and not to comply with the objectives.


EVA 9 Report

Audrey Lan San Reporting


EVA #9, 15:30 - 16:10


Crew members : Audrey Lan San and Romain Benchenafi.
Site location :

Point Papa : N 38.42075 W 110.78280
Point Romeo : N 38.42441 W 110.78572
Refill station : N 38.41784 W 110.78075


Emergency location : N 38.41465 W 110.78056


Transit mode : ATV


Objective : Monitor factors of stress during an EVA


Result : Both targeted samples were found and we intervened on the emergency.


Lesson learned : Always be sure that the oxygen supply is sufficient to fulfill the mission and be back at the Hab. (We almost ran out of oxygen)


Chef's Report

Kévin Théatre Reporting

1a) Today was

-  a cooking day

1b) Was today a special day celebrated at the main meal with special food or activity? No

2)  List the foods served at today's main meal (usually the evening meal), giving full names of each.

Tomatoes, tomato dices, sautéed peppers, carrots, onions.

couscous

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)  Nothing

4)  Recipes for anything more complicated than following package directions exactly, or rehydrating.  Nothing

5)  Lessons learned  (if any) Nothing

6) Comments/questions for Kim and Jean, the food study investigators

We have run out of sliced potatoes and meat.


Journalist Report

Romain Benchenafi Reporting

The first thing we noticed today is the complete lack of wind.

It has been a busy day for all, so everybody is satisfied or  almost satisfied with the work done. We expect to keep that pace until the end of our mission, which, according to the schedule, will come soon.